OpenStudy (anonymous):

how do I solve to find the solution -14< or equal to 3x -7< or equal to -2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[-14\le 3x-7 \le-2 \implies -7\le3x \le5 \implies -{7 \over3} \le x \le {5 \over 3}\]

OpenStudy (anonymous):

AnwarA you are a fast typer XD

OpenStudy (anonymous):

You can use interval notation and write \(x \in [ -\frac{7}{3}, \frac{5}{3}]\).

OpenStudy (anonymous):

@behee: sorry for that :D

OpenStudy (amistre64):

-14 </ 3x -7 </ -2 ; +7 -7 </ 3x </ 5 ; /3 -7/3 </ x </ 5/3 ; Anwar was right yay!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you explain why is it in fraction anwarA

OpenStudy (anonymous):

Because I divide both sides by 3.

OpenStudy (anonymous):

All sides*

OpenStudy (anonymous):

thank you anwarA

OpenStudy (anonymous):

You're welcome. It's just Anwar :D